OpenStudy (anonymous):

HELP WILL FAN AND MEDAL Solve each equation using factoring 16y^3 = 25y

OpenStudy (academicgurusinc):

So first step, you want to take everything to the left side of the equation and set it equal to zero in order to solve the equation. So you'll have 16y^3-25y=0 Then, you'll start factoring. You'll ask yourself what is the most common term here? It is "y" so you'll take out a y and you'll have y(4y-5)(4y+5)=0 You'll solve y(4y-5)(4y+5)=0 and should get y=0 and y = +5/4 and -5/4 Hope that helps. If you found this helpful, I encourage to subscribe to our youtube channel ( https://www.youtube.com/channel/UCYiI7SmkU4_vhdSzKBWsifg), to stay current with all of our new videos. Regards, Academic Gurus Inc. Twitter (@Academic_Gurus) Facebook (AcademicGurusInc) Youtube (Academic Gurus Inc)

OpenStudy (anonymous):

Thanks for helping me.

OpenStudy (academicgurusinc):

You're welcome :)